The length of a lawn is 4 ft more than its width. If the perimeter of the lawn is 40 ft, what is the area of the lawn?
Answer by checkley77(12844)

P=2L+2W
L=W+4
40=2(W+4)+2W
40=2W+8+2W
40=4W+8
4W=40-8
4W=32
W=32/4
W=8 ANS..
L=8+4=12 ANS.
PROOF:
40=2*12+2*8
40=24+16
40=40